# MHNet: Multi-view High-order Network for Diagnosing Neurodevelopmental Disorders Using Resting-state fMRI

### v1

@@ Abstract
Keywords Multi-view; High-order; Neurodevelopmental Disorder; Graph Neural Network; Convolution Neural Network; rs-fMRI; Euclidean Space; Non-Euclidean Space.
@@ 1 Introduction
Autism spectrum disorder (ASD) and attention deficit hyperactivity disorder (ADHD) are two typical neurodevelopmental disorders (NDD) facing specific or sometimes overlapping challenges. Individuals with ASD often exhibit early-onset difficulty in communication and reciprocal social interactions alongside the repetitive and restricted sensory-motor behaviors [1, 2]. On the other hand, individuals with ADHD often show persistent patterns of inattention, hyperactivity, and impulsivity that interfere with daily functioning or development [3, 4].
@@
Currently, the majority of clinical diagnoses of NDD mainly rely on the subjective assessment of abnormal behavior by clinical experts, which may lead to limited accessibility and inconsistent or delayed diagnosis. This highlights the need for more objective computer aided imaging diagnostic methods [7]. As a non-invasive imaging modality, functional magnetic resonance imaging (fMRI) has been adopted for the diagnosis of NDD in previous research studies [5, 6]. Decoding fMRI data using machine or deep learning can provide valuable insights into the abnormality related to NDD including altered neural activity and brain connectivity that are associated with cognitive and executive function.[8], by which the disease-specific changes within the brain can be revealed [9].
@@
Abnormalities or dysfunctions in large-scale brain functional networks (BFNs) can be reflected by altered functional connectivity (FC) derived from resting-state fMRI (rs-fMRI) data. [10]. Numerous studies have discovered that brain functional impairment of NDD patients is related to abnormal FC between resting-state networks (RSNs) [11, 12, 13, 14]. For instance, some studies have found significant differences in the connectivity between the default mode network (DMN) and other brain regions in patients with NDD, which may be associated with the deficits in social behavior and cognitive functions [6, 15]. However, most existing studies only focus on whole brain functional network (BFN) analysis and do not consider the hierarchical structure of the spatial topology of BFNs. This whole BFN analysis does not fully capture the complex changes in brain function of NDD patients. BFNs exhibit significant abnormalities not only in whole BFN but also in local sub-networks, which require multi-scale analytical approaches [16] for constructing networks with hierarchical structure.
@@
Encoding both the Euclidean and non-Euclidean space features based on deep learning can reveal the complementary complex information of the BFNs. The convolutional neural networks (CNNs) can automatically extract the Euclidean space features of BFNs [6, 17, 18, 19], which requires the involvement of local receptive fields encoding the values of connectivity weights of FC. The graph neural networks (GNNs) [20] has emerged as an attractive framework for modeling non-Euclidean space features of BFNs due to their powerful topological graph embedding capabilities [21]. The integration of CNNs and GNNs allows for comprehensive multi-scale feature extraction, enhancing the ability of deep learning model to generalize across different individuals and improve the NDD prediction performance.
@@
However, simply applying CNNs or GNNs only utilizes the first-order features and cannot capture the high-order information in BFNs [22, 23], which limits the ability of generalization and the accuracy of NDD prediction of the deep learning models. In contrast, high-order features are crucial for revealing the complex relationships and indirect connections of the BFNs [24]. Capturing the high-order information of BFNs can potentially improve the generalization ability of the model.
@@ 2.1 Traditional Machine Learning-based NDD Diagnosis
Traditional machine learning methods have been extensively applied in the diagnosis of NDD using fMRI data. These methods typically involve manual feature extraction, feature selection, and classification. Commonly used algorithms include support vector machine (SVM), random forests (RFs), and k-Nearest neighbors (kNN) [30]. SVM has been widely used due to its effectiveness in binary classification tasks such as distinguishing the brain patterns of NDD and health control [31, 32]. RFs and kNN offer robustness and flexibility in handling noisy and complex data. They are employed to classify ADHD patients using rs-fMRI data, showing promising accuracy and interpretability [33, 34]. Traditional machine learning methods heavily depend on the quality of feature extraction and selection, require domain expertise, and is relatively time consuming. Advances like recursive feature elimination and principal component analysis have improved model performance, but the complexity and heterogeneity of NDD still pose challenges, necessitating more advanced techniques [35].
@@ 2.2 GNN-based NDD Diagnosis
GNNs are powerful tools for modeling the topology of BFNs by capturing the complex interactions between brain regions [36, 37, 38]. For instance, park et al. [36] proposed a deep learning model that utilizes a residual graph convolution network (GCN) with spatio-temporal features extracted from 4D fMRI to improve the classification accuracy of ASD. This study focuses on the dynamic FC between superior temporal sulcus and visual cortex. Integrating multiple layers of GCNs can learn hierarchical representations of brain networks characterizing the multi-scale brain dynamics, by which local and global connectivity patterns can be effectively differentiated between ASD and healthy control [37]. Jiang et al. [38] introduced a hierarchical GCN for learning graph embeddings from brain networks to predict brain disorders. This study uses rs-fMRI data to construct brain networks as graphs and enhances the prediction accuracy through hierarchical representation learning. A significant advantage of GNNs is their ability to provide interpretable models. By analyzing the learned node features and edge weights, researchers can gain insights into the encoded information of specific brain regions and connections. Li et al. [39] proposed the BrainGNN which is an interpretable GNN framework to analyze fMRI and discover neurological biomarkers. This method incorporates novel ROI-aware graph convolutional layers and ROI-selection pooling layers to improve the prediction accuracy and interpretability for neurological disorder diagnosis. In addition, GNNs are capable of encoding nuanced alterations of brain connectivity related to NDD according to the study from [40]. Yang et al. [41] proposed a new method named Pearson’s correlation-based spatial constraints representation to estimate the BFNs. In this method, the BFNs were fed into a graph attention network for ASD diagnosis. The development of advanced GNN architectures continues to enhance their applicability and performance, making them a promising tool in the field of NDD diagnosis.
@@ 2.3 CNN-based NDD Diagnosis
CNNs are particularly effective in capturing fine-grained spatial patterns in Euclidean space with local receptive fields [42]. Kawaharaa et al. [18] proposed the BrainNetCNN to predict NDD from structural brain networks of infants, and introduced a special structure with edge-to-edge, edge-to-node, and node-to-graph convolutional layers which can exploit the topological locality of brain networks. Silva et al. [43] developed a CNN-based feature extraction method combining seed correlation, local consistency, and low-frequency amplitude scores for ADHD diagnosis. The 3D CNNs have enhanced capability of capturing volumetric information from fMRI data for NDD diagnosis [44]. Mao et al. [6] designed a variety of spatiotemporal granularity computing and fusion models, including feature pooling, LSTM, and spatiotemporal convolution to capture the spatiotemporal correlation of rs-fMRI. Although challenges remain, CNNs continue to be a promising tool for advancing the field of neuroimaging based NDD diagnosis [45].
@@ 3 Method
In this section, we elucidate the architecture of the proposed MHNet for NDD Diagnosis. Figure 1 shows the MHNet framework which consists of three large modules, and each modules containing several smaller modules: (i) the ESFE module including the components of FCG module and HCNN module; the non-ESFE module including generic G-IBHN-G and three parallel HGNN modules; and FFC module fusing the Euclidean and non-Euclidean features for NDD prediction.
@@ 3.2.2 HGNN
Shallow GNNs have limitation in representing complex network and long-range dependencies, as they typically only capture local node features and neighborhood information [47]. This limitation hinders performance on tasks with diverse FC information for brain disease classification. Although deeper GNNs better capture intricate network structures, they often suffer from over-smoothing, resulting in indistinguishable node features and degrading model performance [48, 49]. The combination of Chebyshev filters and residual connections in ChebNet helps mitigate the over-smoothing problem, which preserves part of original input information at each layer and maintains the diversity of node features across layers [50]. This strategy ensures that the model retains unique characteristic of each node while still benefits from the depth of network.
@@
We define the hierarchically structured brain graphs as $G_{WAN}=(V_{w},A_{w})$, $G_{MAN}=(V_{m},A_{m})$, $G_{LAN}=(V_{l},A_{l})$ where each node represents a sub-graph (or region at the lowest level analysis in G-IBHN-G module) of the corresponding network, i.e. $V_{w}=(v_{1},\ldots,v_{n})$ The node features in $G_{WAN}=(V_{w},A_{w})$ can be expressed as the matrix $H_{w}=[x_{1},x_{2},\ldots,x_{n}]$, where $x_{n}$ is feature expression associated with $v_{n}$.
@@ 3.3 ESFE
The MLP outputs the features $Z_{fc}$, where the transformation of each layer can be expressed as:
@@
where $Z_{HOP}$ is a real symmetric matrix. The expression of the final high-order features is:
@@ 3.4 FFC Module
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right],$ (16)
@@ 4.2 Experimental Setting
The experiments were conducted by using a single server with NVIDIA RTX 3090 GPU. We developed our model using PyTorch, and every algorithm mentioned in this study is capable of running on a single GPU. Adaptive moment estimation (Adam) was employed for network optimization. For ABIDE-I (NC vs. ASD), the dropout rate is set to 0.3, the learning rate is set to 0.0001 and the maximum number of epochs is set to 240, the cutoff threshold is set to 19.03%. For ABIDE-II (NC vs. ASD), the dropout rate is set to 0.25, the learning rate is set to 0.0001 and the maximum number of epochs is set to 200, the cutoff threshold is set to 17.14%. For ADHD-200 (NC vs. ADHD), the dropout rate is set to 0.3, the learning rate is set to 0.0001 and the maximum number of epochs is set to 300, the cutoff threshold is set to 10.23%. The results of classification are averaged over 10 times of cross validated test. Five metrics of AUC, ACC, SEN, SPEC, and AVG are used to evaluate the classification performance.

### v2

@@ Abstract
keywords:
Multi-view; High-order; Neurodevelopmental Disorder; Graph Neural Network; Convolution Neural Network; rs-fMRI; Euclidean Space; Non-Euclidean Space.
@@ 1 Introduction
Autism spectrum disorder (ASD) and attention deficit hyperactivity disorder (ADHD) are two typical neurodevelopmental disorders (NDD) facing specific or sometimes overlapping challenges. Individuals with ASD often exhibit early-onset difficulty in communication and reciprocal social interactions alongside the repetitive and restricted sensory-motor behaviors [1]. On the other hand, individuals with ADHD often show persistent patterns of inattention, hyperactivity, and impulsivity that interfere with daily functioning or development [2].
@@
Currently, the majority of clinical diagnoses of NDD mainly rely on the subjective assessment of abnormal behavior by clinical experts, which may lead to limited accessibility and inconsistent or delayed diagnosis. This highlights the need for more objective computer aided imaging diagnostic methods [3]. As a non-invasive imaging modality, functional magnetic resonance imaging (fMRI) has been adopted for the diagnosis of NDD in previous research studies [4]. Decoding fMRI data using machine or deep learning can provide valuable insights into the abnormality related to NDD including altered neural activity and brain connectivity, by which the disease-specific changes within the brain can be revealed [5, 6].
@@
Abnormalities or dysfunctions in large-scale brain functional networks (BFNs) can be reflected by altered functional connectivity (FC) derived from resting-state fMRI (rs-fMRI) data [7]. Numerous studies have discovered that brain functional impairment of NDD patients is related to abnormal FC between resting-state networks (RSNs) [8, 9]. For instance, some studies have found significant differences in the connectivity between the default mode network (DMN) and other brain regions in patients with NDD, which may be associated with the deficits in social behavior and cognitive functions [4]. However, most existing studies only focus on whole brain functional network (BFN) analysis and do not consider the hierarchical structure of the spatial topology of BFNs. This whole BFN analysis does not fully capture the complex changes in brain function of NDD patients. BFNs exhibit significant abnormalities not only in whole BFN but also in local sub-networks, which require multi-scale analytical approaches [10] for constructing networks with hierarchical structure.
@@
Encoding both the Euclidean and non-Euclidean space features based on deep learning can reveal the complementary complex information of the BFNs. The convolutional neural networks (CNNs) can automatically extract the Euclidean space features of BFNs [4, 11], which requires the involvement of local receptive fields encoding the values of connectivity weights of FC. The graph neural networks (GNNs) [12] has emerged as an attractive framework for modeling non-Euclidean space features of BFNs due to their powerful topological graph embedding capabilities [13]. The integration of CNNs and GNNs allows for comprehensive multi-scale feature extraction, enhancing the ability of deep learning model to generalize across different individuals and improve the NDD prediction performance.
@@
However, simply applying CNNs or GNNs only utilizes the first-order features and cannot capture the high-order information in BFNs [14, 15], which limits the ability of generalization and the accuracy of NDD prediction of the deep learning models. The first-order features are represented by embeddings of individual brain regions extracted from fMRI. In contrast, high-order features are crucial for revealing the complex relationships and indirect connections of the BFNs, encapsulating the intricate correlational patterns among pairwise brain regions, elucidating the topological architecture of neural networks while capturing the multifaceted interactions and emergent global properties inherent in the brain’s organizational framework [16]. Capturing the high-order information of BFNs can potentially improve the generalization ability of the model.
@@ 2.1 Traditional Machine Learning-based NDD Diagnosis
Traditional machine learning methods have been extensively applied in the diagnosis of NDD using fMRI data. Commonly used algorithms include support vector machine (SVM), random forests (RFs), and k-Nearest neighbors (kNN) [22]. SVM has been widely used due to its effectiveness in binary classification tasks such as distinguishing the brain patterns of NDD and health control [23]. RFs and kNN offer robustness and flexibility in handling noisy and complex data. They are employed to classify ADHD patients using rs-fMRI data, showing promising accuracy and interpretability [24, 25]. However, these traditional methods often face computational inefficiencies and limited model generalization capabilities when dealing with high-dimensional and complex fMRI data [4]. Despite improvements from techniques like recursive feature elimination and principal component analysis [26], traditional machine learning methods remain heavily dependent on feature engineering quality and domain expertise, while being time-intensive when handling the complex heterogeneous data of NDD. These limitations necessitate the development of more advanced techniques, such as GNNs and CNNs, which offer enhanced feature learning capabilities and better classification performance to address the shortcomings of traditional methods [27].
@@ 2.2 GNN-based NDD Diagnosis
GNNs are powerful tools for modeling the topology of BFNs by capturing the complex interactions between brain regions. For instance, park et al. [28] proposed a deep learning model that utilizes a residual graph convolution network (GCN) with spatio-temporal features extracted from 4D fMRI to improve the classification accuracy of ASD. Jiang et al. [27] introduced a hierarchical GCN for learning graph embeddings from brain networks to predict brain disorders. A significant advantage of GNNs is their ability to provide interpretable models. By analyzing the learned node features and edge weights, researchers can gain insights into the encoded information of specific brain regions and connections. Li et al. [29] proposed the BrainGNN which is an interpretable GNN framework to analyze fMRI and discover neurological biomarkers. In addition, GNNs are capable of encoding nuanced alterations of brain connectivity related to NDD according to the study from [30]. Yang et al. [31] proposed a new method named Pearson’s correlation-based spatial constraints representation to estimate the BFNs. The development of advanced GNN architectures continues to enhance their applicability and performance, making them a promising tool in the field of NDD diagnosis. While initial GNN applications focused on single modality analysis, recent advances have expanded to leverage multimodal data. The self-attention multi-modal triplet networks have been proposed to synergistically integrate Diffusion Tensor Imaging (DTI) and fMRI modalities to enhance diagnostic precision [32]. HC-GNN employs hypercomplex neural architectures to facilitate multi-modal representation fusion while preserving model interpretability [33]. Furthermore, FC-HGNN implements a dual-stage heterogeneous graph neural framework that incorporates functional connectivity and phenotypic features to facilitate robust biomarker identification [34].
Recent studies have explored advanced GNNs architectures for fMRI analysis, introducing the spatio-temporal hybrid attentive graph network (ST-HAG) [35] that leverages Transformer-based attention for temporal modeling, followed by the explainable spatio-temporal graph convolutional network (ESTGCN) [36] for enhanced interpretability in brain network analysis. Further developments include the explainability-generalizable graph neural network (XG-GNN) [37], which employs meta-learning to improve model generalization, and the causality-inspired graph neural network (CI-GNN) [38] that focuses on identifying diagnostic-relevant subgraphs. However, significant challenges remain as current models struggle with capturing hierarchical features from multi-view brain functional networks and suffer from over-smoothing issues when deepening network layers [39]. Additionally, the limited representation ability of shallow GNN structures and inadequate integration of Euclidean and non-Euclidean features from multi-view data restrict their capacity to fully capture dynamic brain connectivity [40]. Therefore, we proposed a multi-scale residual GNN module within Non-ESFE that captures hierarchical and contextual graph information while addressing limitations in multi-view brain functional networks processing, over-smoothing prevention, and representation capacity enhancement for robust dynamic brain connectivity analysis.
@@ 2.3 CNN-based NDD Diagnosis
CNNs are particularly effective in capturing fine-grained spatial patterns in Euclidean space with local receptive fields [41]. Kawaharaa et al. [11] proposed the BrainNetCNN to predict NDD from structural brain networks of infants, and introduced a special structure with edge-to-edge, edge-to-node, and node-to-graph convolutional layers which can exploit the topological locality of brain networks. Silva et al. [42] developed a CNN-based feature extraction method combining seed correlation, local consistency, and low-frequency amplitude scores for ADHD diagnosis. Multimodal CNN-based approaches have shown promise in neuroimaging analysis. For instance, the integration of multiple MRI modalities through CNN architectures has demonstrated remarkable potential in brain disorder diagnosis [43]. Additionally, Mao et al. [4] designed a 4D CNN architecture that incorporates multiple granularity computing and fusion models, including feature pooling, LSTM, and spatiotemporal convolution to extract dynamic features from rs-fMRI time series. Similarly, Zou et al. [44] proposed a multi-modality 3D CNN architecture that combines fMRI and sMRI features for ADHD diagnosis, highlighting the advantages of brain functional and structural information fusion.
Recent CNN-based models for neuroimaging analysis, such as local-global coupled network (LCGNet) [45] with its dual CNN-transformer backbone and static-dynamic convolutional neural network with dual pathways, have made progress in combining local and global feature extraction [46]. However, these models’ inherent limitation lies in their local receptive fields, which constrain their ability to capture complex, non-Euclidean spatial information and global connectivity patterns of brain networks that extend beyond local neighborhoods. To overcome these limitations, our approach integrates CNNs with GNNs, effectively combining the strengths of both models. This fusion allows us to capture the global topological structures within the brain’s network, providing a more comprehensive understanding of neural dynamics. By leveraging the non-Euclidean capabilities of GNNs alongside the local feature extraction of CNNs, our model addresses the gaps in prior studies.
@@ 3.2.2 HGNN
Shallow GNNs have limitation in representing complex network and long-range dependencies, as they typically only capture local node features and neighborhood information [48]. This limitation hinders performance on tasks with diverse FC information for brain disease classification. Although deeper GNNs better capture intricate network structures, they often suffer from over-smoothing, resulting in indistinguishable node features and degrading model performance [49]. The combination of Chebyshev filters and residual connections in ChebNet helps mitigate the over-smoothing problem, which preserves part of original input information at each layer and maintains the diversity of node features across layers [50]. This strategy ensures that the model retains unique characteristic of each node while still benefits from the depth of network.
The higher-order features can encode more complex relationships and interactions between ROIs. Certain complex interactions in the brain can be easily captured by higher-order correlations, and can exhibit more complex relationships that cannot be directly modeled in lower-order statistics [51]. In the context of GNNs, high-order features extend beyond the immediate neighborhood relationships to capture sophisticated structural and semantic patterns in graph data [52]. To extract these high-order features, we employ a transpose operation that computes pairwise interactions between node embeddings in the latent space. Through this operation, each element in the resulting matrix represents the relationship between a pair of nodes, thereby encoding comprehensive node-to-node interactions across the entire graph [21]. This high-order representation preserves the structural information learned by previous GNN layers while simultaneously capturing more complex topological patterns that may not be directly observable in the original graph structure.
@@
We define the hierarchically structured brain graphs as $G_{WAN}=(V_{w},A_{w})$, $G_{MAN}=(V_{m},A_{m})$, $G_{LAN}=(V_{l},A_{l})$ where each node represents a sub-graph (or region at the lowest level analysis in G-IBHN-G module) of the corresponding network, i.e. $V_{w}=(v_{1},\ldots,v_{n})$. In the construction of these graphs, each node encapsulates specific brain regions or functional units: nodes in $G_{WAN}$ represent major brain systems (e.g., DMN, LN), nodes in $G_{MAN}$ represent intermediate-level brain regions (e.g., FL, TL), and nodes in $G_{LAN}$ represent local brain regions defined by anatomical parcellation. The node features in $G_{WAN}=(V_{w},A_{w})$ can be expressed as the matrix $H_{w}=[x_{1},x_{2},\ldots,x_{n}]$, where $x_{n}$ is feature expression associated with $v_{n}$. The edges in these graphs, represented by adjacency matrices $A_{w}$, $A_{m}$, and $A_{l}$, are constructed based on both structural and functional connectivity. The edge weights $w_{ij}$ between nodes $i$ and $j$ are determined by:
$A_{k}(i,j)=\begin{cases}RV_{ij}&\text{if }\rho_{ij}>\gamma\\ 1&\text{if }i=j\\ 0&\text{otherwise},\end{cases}$ (4)
where $k\in{w,m,l}$ represents the network level, $\rho_{ij}$ is the RV correlation coefficient between the BOLD signals of regions $i$ and $j$, and $\gamma$ is a threshold parameter to control the graph sparsity.
For graphs $G_{MAN}$ and $G_{LAN}$, we partition the brain regions into subgraphs based on functional networks or structural partitions. These subgraphs are arranged along the diagonal of a larger sparse adjacency matrix, forming a block-diagonal matrix:
$A_{k}=\operatorname{diag}\left(A_{1},A_{2},\ldots,A_{n}\right),$ (5)
where each $A_{i}$ is the adjacency matrix of the $i$-th subgraph, $k\in{m,l}$. This configuration enables us to process all subgraphs collectively through GNN layers while preserving their individual topological characteristics.
@@ 3.3 ESFE
Traditional CNN layers recognize local patterns through convolution, but high-order representations model complex feature interactions in deeper layers, capturing abstract patterns beyond local feature detection. Deeper layers learn abstract and composite features by combining lower-level elements, leading to more discriminative representations [54]. High-order pooling operations capture correlations and dependencies that conventional linear operations in CNN may miss, enriching feature representations that are particularly important in medical image analysis, where complex patterns and subtle changes are crucial for accurate diagnosis [55].
@@
The MLP outputs the features $Z_{fc}\in\mathbb{R}^{d}$, where the transformation of each layer can be expressed as:
@@
where $Z_{HOP}$ is a real symmetric matrix. Each element $Z_{HOP}(i,j)$ in the matrix can be expressed as: $Z_{HOP}(i,j)=\langle Z_{fc}(i,:),Z_{fc}(j,:)\rangle$, which reflect the interaction strength and covariation relationship between them. This operation captures pairwise multiplicative interactions between different feature dimensions [56], enabling the extraction of interactions and co-variations between functional connectivity patterns, and providing richer high-order statistical features beyond the original FC.
The expression of the final high-order features is:
@@ 3.4 FFC Module
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$ (18)
@@ 4.2 Experimental Setting
Figure 2: Training loss curves on three datasets. The plots show the convergence of the model during training: (a) ABIDE-I dataset with 240 epochs, (b) ABIDE-II dataset with 200 epochs, and (c) ADHD-200 dataset with 300 epochs.
The experiments were conducted by using a single server with NVIDIA RTX 3090 GPU. We developed our model using PyTorch, and every algorithm mentioned in this study is capable of running on a single GPU. Adaptive moment estimation (Adam) was employed for network optimization. For ABIDE-I (NC vs. ASD), the dropout rate is set to 0.3, the learning rate is set to 0.0001 and the maximum number of epochs is set to 240, the cutoff threshold is set to 19.03%. For ABIDE-II (NC vs. ASD), the dropout rate is set to 0.25, the learning rate is set to 0.0001 and the maximum number of epochs is set to 200, the cutoff threshold is set to 17.14%. For ADHD-200 (NC vs. ADHD), the dropout rate is set to 0.3, the learning rate is set to 0.0001 and the maximum number of epochs is set to 300, the cutoff threshold is set to 10.23%. The results of classification are averaged over 10 times of cross validated test, where a train/validation/test ratio of 70%/10%/20% was consistently applied across all comparison methods with balanced class distribution. The training convergence curves of our model on three datasets are shown in Fig. 2. The loss curves demonstrate stable and consistent convergence patterns across all datasets, with rapid initial descent in the early epochs followed by gradual stabilization, indicating effective learning and parameter optimization. Five metrics of AUC, ACC, SEN, SPEC, and AVG are used to evaluate the classification performance.
@@ 4.4 Evaluation Metrics
To compare the performance of models, we use the almost stochastic order (ASO) test of statistical significance. For all tests, the significance level is set to $\alpha$ = 0.05 and adjusted using the Bonferroni correction when making multiple comparisons.